The index of refraction for violet light in silica flint glass is 1.66 and that for red light is 1.62. What is the angular dispersion of visible light passing through an equilateral prism of apex angle 60.0
Answer #1

For the violet light, 1.66, and nar = 1.00 600Y 1.00sin 56 θ2r -sin- 29.96 1.66 α-90.-6,ー60.0, β-180.0°-60.0°-α-5996 2slassangular dispersion of visible light is 56.197-52.26= 3.935 deg---> Answer
Answer #2

The index of refraction for violet light in silica flint glass is 1.66, and that for red light is 1.62. What is the angular dispersion of visible light passing through a prism of apex angle 60.0° if the angle of incidence is 50.0°? (See figure)
Answer #3

for violet glass

n1sin i = n2sin r1

r1 = 28.34 deg

i2 = A - r1= 31.66 deg


n1sin 31.66 = n2sinr2

r2 = 60.60 deg


for red glass

n1sin i = n2sin r1

r1 = 29.11 deg

i2 = A - r1= 30.89 deg


n1sin 30.89 = n2sinr2

r2 = 56.27 deg


angular despersion = 60.6 - 56.27 = 4.33 deg